[Brahms-l] Shift report 20050418 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Mon Apr 18 2005 - 15:50:46 EDT
Shifter leader was BB 
Shifters were JHL


1. Beam for the first 2 hours. Machine development
for rest of the shift.

2.
Run 14079 trig3(31k), trig4(122k)
Run 14080 trig3(35k), trig4(144k)
Run 14081 non-TPC pedestal run
Run 14082 TPC pedestal run

3. No detector problems



_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Apr 18 15:50:59 2005

This archive was generated by hypermail 2.1.8 : Mon Apr 18 2005 - 15:51:15 EDT